Circular No. 8458

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 Y1 (CHRISTENSEN) Eric J. Christensen, Lunar and Planetary Laboratory, reports his dis-

covery of a comet in the course of the Catalina Sky Survey on images taken with the 0.68-m Schmidt telescope (discovery observation below), noting a coma of diameter ∼ 15′′ and a fan-shaped tail 20′′ long in p.a. 270^{o}. Following posting on the ‘NEO Confirmation Page’, J. Young writes that his images taken at Table Mountain on Dec. 22.4 UT show the object with a round 8′′ coma and a broad fan-shaped tail spanning p.a. 210^{o}–290^{o} (the faint tail edge at p.a. 210^{o} extends 20′′, while the brighter edge at p.a. 290^{o} extends only 14′′).

2004 UT  α_{2000}  δ_{2000}  Mag.
Dec. 21.47360  11^{h}02^{m}23.^{s}99  +22^{o}52′57′′.8  16.5

The available astrometry (including prediscovery observations by LINEAR back to Dec. 9), the following orbital elements, and an ephemeris appear on MPEC 2004-Y47.

T = 2004 June 21.588 TT  ω = 325.^{o}364
e = 0.54459  Ω = 88.689 }2000.0
q = 1.64584 AU  i = 16.242
a = 3.61398 AU n^{o} = 0.143458 P = 6.87 years

COMETS C/2004 V14, V15, W1–W6 Further to IAUC 8455, following are the initial observations of addi-

tional Kreutz sungrazing comets found on SOHO website images (TS = T. Scarmato). K. Battams notes that C/2004 V15 was fairly bright with a short tail.

Comet 2004 UT α_{2000} δ_{2000} Inst. F MPEC
C/2004 V14 Nov. 14.096 15^{h}11.^{m}2 −20^{o}43′ C3/2 MM 2004-Y37
C/2004 V15  15.904 15 20.6 −22 20 C3/2 TH 2004-Y37
C/2004 W1 16.254 15 24.9 −20 20 C2 TH 2004-Y37
C/2004 W2  18.513 15 31.9 −21 34 C3/2 TS 2004-Y37
C/2004 W3  19.513 15 35.6 −22 38 C3/2 RK 2004-Y38
C/2004 W4  21.785 15 49.0 −21 45 C2  BZ 2004-Y38
C/2004 W5  22.368 15 52.1 −21 51 C2  SH 2004-Y38
C/2004 W6  23.213 15 55.6 −22 05 C2  XL 2004-Y38
2004 December 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green